calculate the volume of air needed for teh combustion of 1 kg of carbon?
 

Grade:10

1 Answers

swarup kumar majee
6 years ago
C+O2=CO2
SO atomic mass of C=12 g     O2=32g
32g O2required to combuss   12g C
SO 32/12*1000=8000/3=266.6g  O2 required to combuss 1kg C